# Design and Evaluation of SEANet: a Software-defined Networking Platform for the Internet of Underwater Things

### v1

@@ VI Conclusions
The paper concerns the design, prototyping, and testing the SEANet software-defined networking platform for underwater wireless networks. The hardware and software design principles guiding the realization of the platform are explained, along with its prototyping and related challenges. Results from experiments at sea performed by mounting an acoustic front-end showcase the platform capabilities and its ease of use. Thanks to its software-defined architecture, in the same experimental run, the SEANet networking platform can perform channel and frequency response, establish an OFDM communication link with different bandwidths and modulations, achieve more than double the data rates of the fastest commercial modems, can establish a bidirectional ARQ communication link, and can communicate using JANUS, the communication standard for underwater communications.

### v2

@@ VI Conclusions
The paper concerns the design, prototyping, and testing of the SEANet software-defined networking platform for underwater wireless networks. The platform has been engineered to address the prevalent challenges in underwater communication by integrating flexibility, high data throughput, and real-time adaptability into its core design. Throughout various deployments and rigorous testing in oceanic environments, SEANet has demonstrated its effectiveness by significantly outperforming existing commercial underwater modems in terms of data rate, achieving more than double the data rates of the fastest commercial modems. Our experimental results have validated the platform versatility across multiple aspects, including its ability to perform reliable data transmission, establish OFDM links with different bandwidths and modulations, and integrate seamlessly with standard communication protocols like JANUS. Furthermore, the adaptability of SEANet is highlighted through its successful operation under different configurations and conditions, showcasing its potential to support diverse underwater operations—from data-heavy scientific research to critical infrastructure monitoring. In conclusion, the SEANet platform stands as a significant advancement in the field of underwater communication, offering a scalable, flexible, and high-performance solution with the potential to revolutionize how we deploy and utilize the Internet of Underwater Things.